OpenStudy (anonymous):

Ann claims that she can't show 3/4 of the following faces b/c some are big and some are small. What do you tell her? FYI there are a total of 8 faces

OpenStudy (anonymous):

clarify....

OpenStudy (anonymous):

Clarify the different faces I see? Is that what you are asking?

OpenStudy (anonymous):

There are 4 big faces, 4 small faces, and of the 4 small faces 1 has a frown

hero (hero):

lol

hero (hero):

Well, in that case, what you want to do is set up a proportion: \(\frac{3}{4} = \frac{x}{8}\) Now solve for x: 3(8) = 4x 24 = 4x 24/4 = 6 So now you have the ratio 6/8 which is equivalent to 3/4. Thus 3 of the big figures + 3 of the smaller figures represents 3/4 of the figure. That explanation may or may not make any sense to you. But anyway, since 6/8 = 3/4, 6 faces out of eight still represents 3/4 of the figure.

OpenStudy (anonymous):

I thank you so much I was setting it up like 3/4 * 8/1 but I kept thinkin something is not right what is going on here but I will rewrite it to understand it, I am a hands-on learner so doing this math online can be challenging at times for me but thanks again for the breakdown